OpenStudy (pottersheep):

Grade 11 Molecular Mass Help please

OpenStudy (pottersheep):

Calculate the molecular weight of a gas if 4.5L of the gas at 104KPA at 296.5K weighs 13.5g. The catch is I HAVE to use the forula v1p1/t2 = v2p2/t2 and NOT the ideal gas law Please help me I don’t know how finding the VOLUME willhelp me find the molecular weiht..

sam (.sam.):

PV=nRT

OpenStudy (pottersheep):

can't:( i cant use the ideal gas law

sam (.sam.):

PV=nRT \[PV=\frac{13.5g}{M_r}RT\] \[M_r=\frac{13.5g}{PV}RT\] \[M_r=\frac{13.5g}{(1.026atm)(4.5L)}(0.082\frac{atmL}{molK})(296.5K)\] \[M_r=71.09g/mol\]

OpenStudy (pottersheep):

thank you :) that's what I did but my teacher told us to find a way to do it with the other formula but i don't even know if it's possible

sam (.sam.):

you can't use v1p1/t2 = v2p2/t2 because you're not comparing gases

OpenStudy (pottersheep):

She said convert them to STP conditions

OpenStudy (pottersheep):

If I do, I find the volume. Then I couldnt find a way to calculate the mass...

OpenStudy (pottersheep):

same here, im glad im not the only one then, maybe she's thinking wrong. Thanks for your help!

OpenStudy (anonymous):

Here's how you use the V1P1/T1 = V2P2/T2 method. At given conditions: 1.027 atm 296.5 K 4.5 L 13.5 g At STP conditions: 1 atm 273 K ? L 13.5 g (4.5 L)(1.027 atm)/296.5 K = (V2)(1 atm)/273 K V2 = 4.26 L Since V2 is volume at STP, you may convert it to moles pretty easily with this conversion factor: 1 mol = 22.4 L 0.19 mol 13.5g/0.19 mol = 71 g/mol

OpenStudy (pottersheep):

Oh myy thanks sooo much! :D haha no wonder